solve for x and y using elimination method 
x+y/xy=2
x-y/xy=6

Grade:10

1 Answers

Shubh Baghel
6 years ago
Hi, you may find many different solutions but this is the simplest. ... but check the answer first and then follow my steps... x+y/xy =2 x-y/xy =6x+y=2xy.......... (1) x-y=6xy ....... (2)Add 1 and 22x=8xy2=8yY=1/4